# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

### v1

@@ IV.2 Dichroism of the Optical Circular Conductivity
for the Right-handed ($\sigma_{+}$) and Left-handed ($\sigma_{-}$) optical circular conductivity. It is important to note here that we work under the assumption that the monoclinic tilt is non-negligible and therefore take the orthorhombic approximation because the optic axes in our measurement stay nearly in-plane. We also note that $\sigma_{xy}$ should be independent of the actual axis in which we have resolved the Faraday angle rotation. Moreover, while linear birefringence mixes the two circular eigenmodes, so they are not perfectly circular, it does not affect the Faraday angle rotation since it is not related to the off-diagonal terms and cancels out in the cross-polarization measurement. However, the values of $\sigma_{xy}$ can be different for the measurement and analysis of the Faraday angle rotation in Pol1 vs Pol2 due to systematic errors, e.g. residual cross polarization signal, and/or crystal misalignment. Our results show that throughout most of the spectral range, including the lower energy magnon mode and phonons, there is a perfect agreement between the two sets of data with respect to $\sigma_{xy}$ obtained from the two in-plane polarizations. Some discrepancy between the values can be detected around the upper energy magnon mode (see App. VII), which shows that it is local in frequency and probably not related to systematic errors or any sample tilt. Therefore, we have averaged the two $\sigma_{xy}$ outputs in order to obtain the optical circular conductivity $\sigma_{\pm}$.
@@
Figure 10 shows the circular optical conductivity obtained from the analysis of the Magneto-Optical Kerr/Faraday Effect [23] and by taking into account the in-plane optical conductivity tensor as described above. As expected from a collinear antiferromagnetic resonance, there is a clear dichroism between $\sigma_{+}$ and $\sigma_{-}$ under the application of a magnetic field seen at frequencies of about 116 $cm^{-1}$ and 129 $cm^{-1}$ for 7 T for the lower energy and upper energy, respectively, of the 122 $cm^{-1}$ zero-field magnon.
@@
In fact, Fig.  10 shows that the lower mode is completely dichroic since the spectral weight of $\sigma_{-}$ in that excitation is the same magnitude as that of $\sigma_{1}$ at zero field while $\sigma_{+}$ shows a clear peak. On the other hand, the upper energy mode at around 129 $cm^{-1}$ does not show an ideal dichroic response with $\sigma_{+}$ being about factor of 1/3 from the value of $\sigma_{-}$. This response is not subject to the average of $\sigma_{xy}$ of the two polarizations and can be seen in both sets of data for the upper energy magnon mode.
@@
Since this anomalous dichroism appears only in the upper energy magnon which is also close to an infra-red active phonon, we suggest that spin-phonon coupling is responsible for prohibiting optical transitions between the two circular energy states, thus giving a residual, but still significant, spectral weight to $\sigma_{+}$ which is supposed to be closer to $\sigma_{1}$ at zero field in that area (around $1\leavevmode\nobreak\ S$ $cm^{-1}$). Previous reports of coupling between the lower magnon mode and an optical phonon have been reported as an avoided crossing in the optical spectra [34]. Therefore, we suggest that the upper magnon mode is already developing an avoided transition due to spin-phonon coupling involving the upper energy magnon branch and the infrared active phonon at 133 $cm^{-1}$.
@@ V Conclusions
We have conducted a comprehensive infrared spectroscopic investigation of FePS3. By combining absolute transmission, optical conductivity, and Faraday rotation measurements, we explored its collective excitations. Our measurements reveal that the lattice vibrational modes exhibit marked anisotropy and field sensitivity in the antiferromagnetic state, providing direct evidence of strong spin–lattice coupling. A key feature is the 122 $cm^{-1}$ magnon mode, which appears in both infrared and Raman spectra. It exhibits isotropic temperature evolution and undergoes magnetic-field-induced splitting, consistent with antiferromagnetic order.
Significantly, several phonon modes display pronounced Faraday rotation, and certain Raman-active modes appear in the infrared spectra at zero field—features suggesting inversion symmetry breaking, likely driven by spin–phonon interactions. These observations challenge earlier reports suggesting that inversion symmetry is preserved in FePS3, highlighting the need for further studies to reconcile these findings. Moreover, our analysis of the circular optical conductivity reveals dichroism associated with magnon splitting. Anomalous behavior near certain phonon energies points to avoided crossings and magneto-elastic hybridization.
Our results reveal a rich magneto-optical landscape in FePS3 and demonstrate the critical role of spin–lattice interactions in shaping its collective excitation spectrum. Future experiments combining ultrafast optical probes or monolayer measurements could provide deeper insight into the dynamics of spin–lattice coupling and its potential for controlling magnetic phases in two-dimensional materials, and for their potential applications in spintronic and magneto-optical technologies.
@@ Acknowledgments
VI Transmission Data
@@
In Figure 11 we show a comparison of the transmission data measured for the two principal polarizations at 150 K, which is above the Néel temperature, TN=118 K, and at 5 K, which is deep in the antiferromagnetic state. At 150 K, there is no significant difference between the two spectra, except for the phonon at around 220 $cm^{-1}$, indicating isotropic transmission behavior in the paramagnetic phase and on the $a-b$ plane of the monoclinic structure of FePS3. However, as the temperature approaches TN, noticeable changes appear in the spectra, making it impossible to fit the data with a single isotropic model and therefore pointing to the onset of the magnetic order and increasing fluctuations of the magneto-elastic effects. The same spin fluctuations are also responsible for the change in the mode around 100 $cm^{-1}$seen in the Raman spectrum [30]. Below $T_{N}$, there is an abrupt change in the spectra up to about 250  $cm^{-1}$owing to the Brillouin zone folded phonons and appearance of the magnon mode, as was seen also previously in Raman spectroscopy.
VII Optical Circular Conductivity

### v2

@@ IV.3 Dichroism of the Optical Circular Conductivity
for the Right-handed ($\sigma_{+}$) and Left-handed ($\sigma_{-}$) optical circular conductivity. It is important to note here that we work under the assumption that the monoclinic tilt is non-negligible and therefore take the orthorhombic approximation because the optic axes in our measurement stay nearly in-plane. We also note that $\sigma_{xy}$ should be independent of the actual axis in which we have resolved the Faraday angle rotation. Moreover, while linear birefringence mixes the two circular eigenmodes, so they are not perfectly circular, it does not affect the Faraday angle rotation since it is not related to the off-diagonal terms and cancels out in the cross-polarization measurement. However, the values of $\sigma_{xy}$ can be different for the measurement and analysis of the Faraday angle rotation in Pol1 vs Pol2 due to systematic errors, e.g. residual cross polarization signal, and/or crystal misalignment. Our results show that throughout most of the spectral range, including the lower energy mode (approx. 116 $cm^{-1}$) and phonons, there is a perfect agreement between the two sets of data with respect to $\sigma_{xy}$ obtained from the two in-plane polarizations. Some discrepancy between the values can be detected around the upper energy mode (approx. 129  $cm^{-1}$, see App. B), which shows that it is local in frequency and probably not related to systematic errors or any sample tilt. Therefore, we have averaged the two $\sigma_{xy}$ outputs in order to obtain the optical circular conductivity $\sigma_{\pm}$.
@@
Figure 12 shows the circular optical conductivity obtained from the analysis of the Magneto-Optical Kerr/Faraday Effect [36] and by taking into account the in-plane optical conductivity tensor as described above. As expected from a collinear antiferromagnetic resonance, there is a clear dichroism between $\sigma_{+}$ and $\sigma_{-}$ under the application of a magnetic field seen at frequencies of about 116 $cm^{-1}$ and 129 $cm^{-1}$ for 7 T for the lower energy and upper energy, respectively, of the 122 $cm^{-1}$ zero-field mode.
@@
In fact, Fig.  12 shows that the lower mode is completely dichroic since the spectral weight of $\sigma_{-}$ in that excitation is the same magnitude as that of $\sigma_{1}$ at zero field while $\sigma_{+}$ shows a clear peak. On the other hand, the upper energy mode at around 129 $cm^{-1}$ does not show an ideal dichroic response with $\sigma_{+}$ being about factor of 1/3 from the value of $\sigma_{-}$. This response is not subject to the average of $\sigma_{xy}$ of the two polarizations and can be seen in both sets of data for the upper energy mode (129 $cm^{-1}$).
@@
Since this anomalous dichroism appears only in the upper energy mode (129  $cm^{-1}$) which is also close to an infra-red active phonon, we suggest that spin-phonon coupling is responsible for prohibiting optical transitions between the two circular energy states, thus giving a residual, but still significant, spectral weight to $\sigma_{+}$ which is supposed to be closer to $\sigma_{1}$ at zero field in that area (around $1\leavevmode\nobreak\ S$  $cm^{-1}$). Previous reports of coupling between the lower energy mode (116  $cm^{-1}$) and an optical phonon have been reported as an avoided crossing in the optical spectra [56]. Therefore, we suggest that an avoided transition due to spin-phonon coupling involving the upper energy mode (129  $cm^{-1}$) and the infrared active phonon at 133  $cm^{-1}$is already present at low fields.
@@ V Conclusions
We reported a comprehensive magneto-optical study of FePS3 that resolves the interplay between magnetic and lattice degrees of freedom. In the antiferromagnetic state, the phonon spectrum becomes strongly anisotropic. On the other hand, the $122$  $cm^{-1}$ ($15$ meV) excitation is polarization-independent, hardens with an order-parameter-like temperature dependence, and splits with field at low temperatures, producing a large Faraday rotation. By combining absolute transmission with full-protocol Faraday measurements, we extracted the circular optical conductivities and uncovered a clear dichroism of the split excitations. The asymmetric spectral weight of the upper branch near $129$  $cm^{-1}$ is naturally explained by spin–phonon coupling to a nearby infrared phonon. Sizable Faraday rotation on several phonons, together with the appearance in infrared of Raman-active modes and a broad band near $900$  $cm^{-1}$ below $T_{\mathrm{N}}$, indicates inversion-symmetry breaking in the magnetic phase. A weak Faraday signature around $320$  $cm^{-1}$ ($\sim 40$ meV) points to an additional magnetic excitation. These findings highlight a rich magneto-optical landscape in FePS3, motivate future experiments combining ultrafast optical probes or monolayer extensions, and demonstrate a general route to quantify spin–lattice coupling and circular dichroism in two-dimensional antiferromagnets.
@@ Acknowledgments
Appendix A Transmission Data
As explained in the main text and in Sec. II, FePS3h̃as an anisotropic response in the antiferromagnetic state. In order to identify two principal optical axes for our polarization-dependent measurements, we have measured the transmission spectrum of our sample at 5 K (deep in the AFM state) and at zero field as a function of the incoming polarizer angle. Figure 13 shows a colormap plot of the polarization-dependent transmission spectra in which two distinct spectra are identified as the principal optical axes of the FePS3s̃ingle crystal sample. The same procedure was conducted in the MIR range using a KRS-5 polarizer and associated with their counterpart FIR polarized data. The entire set of transmission data from the FIR and MIR ranges was then combined into two sets termed as "Pol 1" and "Pol 2" in our work.
Figure 13: Colormap plot of the transmission spectra as a function of the incoming polarizer (Gold polarizer for the FIR range) angle in the range of 50 to 250  $cm^{-1}$. The gray dashed lines represent the two principal polarizations $P$ used in this work for the temperature and magnetic field-dependent transmission measurements. Here $P=150^{\circ}$ is defined as "Pol 1" and $P=60^{\circ}$ as "Pol 2". The 122 $cm^{-1}$ excitation can be seen in this colormap as a polarization-independent absorption compared to the lattice collective modes.
@@ Appendix A Transmission Data
In Figure 14 (a)-(b) we show a comparison of the transmission data measured for the two principal polarizations at 150 K, which is above the Néel temperature, TN=118 K, and at 5 K, which is deep in the antiferromagnetic state. At 150 K, there is no significant difference between the two spectra, except for the phonon at around 220 $cm^{-1}$, indicating isotropic transmission behavior in the paramagnetic phase and on the aba-b plane of the monoclinic structure of FePS3. However, as the temperature approaches TN, noticeable changes appear in the spectra, making it impossible to fit the data with a single isotropic model and therefore pointing to the onset of the magnetic order and increasing fluctuations of the magneto-elastic effects. The same spin fluctuations are also responsible for the change in the mode around 100 $cm^{-1}$ seen in the Raman spectrum [43]. Below $T_{N}$, there is an abrupt anisotropy in the spectra up to about 250  $cm^{-1}$ since the three-fold rotational symmetry in FePS3 is broken by the formation of the zigzag spin chains, resulting also in the appearance of Brillouin zone folded phonons as was seen also previously in Raman spectroscopy [15, 14].
To emphasize the effect of the magnetic state on the anisotopic optical response of FePS3  we present in Figure 14(c) the zero-field transmission data for MnPS3 at two temperatures: 100 K ($T>T_{N}$) and 10 K ($T\ll T_{N}$), where $T_{N}=78$ K. The spectra show that in the antiferromagnetic state of MnPS3, there are no new phonon modes. Both MnPS3 and FePS3 have the same monoclinic crystal structure with a C2/m space group. However, their magnetic structure in the antiferromagnetic state is different. In FePS3, the unit cell doubles along one direction due to the zigzag magnetic structure and the formation of ferromagnetic chains along the $a$-axis and antiferromagnetic coupling between the chains along the $b$-axis. As a result, new phonon modes, which are folded from higher momentum points to the $\Gamma$ point, emerge below $T_{N}$ [15, 14, 60, 43]. In contrast, MnPS3 has an isotropic Néel spin configuration, meaning that each magnetic moment is antiparllel to all its nearest neighbors. As a result, there is no need to consider a larger unit cell in order to incorporate this structure, as happens in FePS3. Therefore, no new phonon modes appear in its spectrum, consistent also with the Raman spectroscopy measurement [61].
Figure 14: Transmission spectra of FePS3 ($T_{N}=118$ K) for two polarizations at (a) 150 K ($T>T_{N}$) and (b) 5 K ($T\ll T_{N}$). (c) Zero-field transmission spectra at 100 K ($T>T_{N}$) and 10 K ($T\ll T_{N}$) for MnPS3 ($T_{N}=78$ K). No additional phonon modes emerge in the antiferromagnetic state.
Appendix B Optical Circular Conductivity in the Birefringence Case
We start by defining a dielectric tensor of an anisotropic sample including the off-diagonal magneto optical terms which is given by:
$\epsilon(\omega)=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}&0\\ \epsilon_{yx}&\epsilon_{yy}&0\\ 0&0&\epsilon_{zz}\end{pmatrix}$ (4)
where $x,y,z$ are the cartesian axes. Since the light is propagating in the $z$ direction, we can refer to the in-plane dielectric tensor $e_{\perp}$, which is a $2\times 2$ matrix, in the following derivation. It is important to note that our derivation follows the orthorhombic approximation since we neglect, at the moment, the tilt of the monoclinic structure of FePS3. Therefore, the dielectric tensor in the xyx-y plane:
$\epsilon_{\perp}=\begin{pmatrix}\epsilon_{xx}&\epsilon_{xy}\\ -\epsilon_{xy}&\epsilon_{yy}\end{pmatrix}$ (5)
where the in-plane anisotropy is given by the fact that $\epsilon_{xx}\neq\epsilon_{yy}$ and the off diagonal element is $\epsilon_{yx}=-\epsilon_{xy}$.
In the same way we can define the in-plane optical conductivity tensor:
$\sigma{\perp}=\begin{pmatrix}\sigma_{xx}&\sigma_{xy}\\ -\sigma_{xy}&\sigma_{yy}\end{pmatrix}$ (6)
since the dielectric function and optical conductivity are connected via:
$\epsilon_{\mu\nu}(\omega)=\delta_{\mu\nu}+\frac{i}{\epsilon_{0}\omega}\sigma_{\mu\nu}(\omega)$ (7)
where $\mu,\nu\in{x,y}$, $\delta_{\mu\nu}$ is the Kronecker delta, $\epsilon_{0}$ is the vacuum permittivity, and $\omega$ is the optical frequency.
The circular polarization basis is given by the eigenvectors:
$\hat{e}_{R,L}=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ \mp i\end{pmatrix}$ (8)
where $L$ represents a clockwise rotation or Left-Handed (LH) circular polarization, while $R$ represents a counterclockwise rotation or Right-Handed (RH) circular polarization.
In order to convert from the xyx-y basis to the circular basis, we need to find the matrix $U$ which will convert it as follows:
$\begin{pmatrix}\hat{x}\\ \hat{y}\end{pmatrix}=U\begin{pmatrix}\hat{e}_{R}\\ \hat{e}_{L}\end{pmatrix}$ (9)
which is:
$U=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ i&-i\end{pmatrix}$ (10)
while for the inverse case:
$\begin{pmatrix}\hat{e}_{R}\\ \hat{e}_{L}\end{pmatrix}=U^{\dagger}\begin{pmatrix}\hat{x}\\ \hat{y}\end{pmatrix}$ (11)
we write that:
$U^{\dagger}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ 1&i\end{pmatrix}$ (12)
where $U$ is a unitary matrix and $U^{\dagger}U=UU^{\dagger}=\mathbb{1}$
We can now convert the dielectric tensor from the $xy$ basis to the circular basis as follows:
$\epsilon_{circ}=U^{\dagger}\epsilon_{\perp}U=\begin{pmatrix}\epsilon_{+}&\delta\\ \delta&\epsilon_{-}\end{pmatrix}$ (13)
where the diagonal elements are:
$\epsilon_{\pm}=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}\pm i\epsilon_{xy}$ (14)
and the off-diagonal elements, representing the birefringence, are:
$\delta=\frac{\epsilon_{xx}-\epsilon_{yy}}{2}$ (15)
Here the (++) sign refers to the LH circular dielectric element, $\epsilon_{+}=\epsilon_{LL}=\hat{e}^{\dagger}_{L}\epsilon\hat{e}_{L}$, the (-) sign refers to the RH circular dielectric element, $\epsilon_{-}=\epsilon_{RR}=\hat{e}^{\dagger}_{R}\epsilon\hat{e}_{R}$, and $\delta=\epsilon_{RL}=\epsilon_{LR}$.
If $\delta=0$ or $\epsilon_{xx}=\epsilon_{yy}$ then the eigenmodes of the dielectric tensor are purely circular where $n_{L}^{2}=\epsilon_{+}=\epsilon_{LL}$ for LH circular polarization and $n_{R}^{2}=\epsilon_{-}=\epsilon_{RR}$ for RH circular polarization and $n=\sqrt{\epsilon}$ is the refractive index. Ref. [22] follows this isotropic case in order to extract the Faraday (Kerr) angle rotation from the magneto-transmission (-reflection) data. However, since FePS3 is anisotropic, we have to go through the entire derivation of the Faraday angle in the case of a birefringent sample.
For linear polarized light which is a superposition of LH and RH circular polarization, we can write that:
$\vec{E}_{in}=E_{0}\hat{x}=\frac{E_{0}}{\sqrt{2}}(\hat{\epsilon}_{L}+\hat{\epsilon}_{R})$ (16)
The incoming field propagates through a slab of thickness $d$, where each circular component will accumulate a different phase:
$\vec{E}_{out}=\frac{E_{0}}{\sqrt{2}}(\hat{\epsilon}_{L}e^{ikn_{L}d}+\hat{\epsilon}_{R}e^{ikn_{R}d})$ (17)
The phase difference $\Delta\phi=\frac{\omega d}{c}(n_{L}-n_{R})$ between the two modes will result in the Faraday rotation angle:
$\theta_{F}=\frac{1}{2}\Re{\left[\frac{\omega d}{c}(n_{L}-n_{R})\right]}$ (18)
and the ellipticity:
$\eta_{F}=\frac{1}{2}\Im{\left[\frac{\omega d}{c}(n_{L}-n_{R})\right]}$ (19)
A nonzero $\delta$ implies that the sample has a linear birefringence term and the dielectric tensor in the circular basis is not diagonal, nor $\hat{\epsilon}_{R}$ and $\hat{\epsilon}_{L}$ are the true purely circular eigenmodes. The polarization eigenstates in this case tilt away from the circular axes and mix the LH and RH components.
The eigenvalues of $\epsilon_{circ}$ are therefore obtained by diagonalizing the matrix $\mathrm{det}(\epsilon_{circ}-\lambda_{\pm}\mathbb{1})=0$ but recognize that $n_{\pm}=\sqrt{\lambda_{\pm}}$, Therefore we can write the refractive indices as::
$n^{2}_{\pm}$ $=\frac{\epsilon_{+}+\epsilon_{-}}{2}\pm\sqrt{\left(\frac{\epsilon_{+}-\epsilon_{-}}{2}\right)^{2}+\delta^{2}}$ (20)
$=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}\pm\sqrt{\delta^{2}-\epsilon_{xy}^{2}}$ (21)
where the first term in the right-hand side of Eq. 21 is the average of the in-plane dielectric functions:
$\bar{\epsilon}=\frac{\epsilon_{xx}+\epsilon_{yy}}{2}$ (22)
Below the square root are the linear birefringence term:
$\delta=\frac{\epsilon_{xx}-\epsilon_{yy}}{2}$ (23)
and the off-diagonal term which corresponds to the Magneto-Optical effect on the optical functions of the sample. Therefore we can write Eq. 21 as follows:
$n^{2}_{\pm}=\bar{\epsilon}\pm\sqrt{\delta^{2}-\epsilon_{xy}^{2}}$ (24)
We can define $\epsilon_{xx}=n_{0}^{2}e^{+\theta}$ and $\epsilon_{yy}=n_{0}^{2}e^{-\theta}$ with a common scale $n_{0}^{2}=\sqrt{\epsilon_{xx}\epsilon_{yy}}$ and an exponential factor $\theta=0.5\ln(\epsilon_{xx}/\epsilon_{yy})$ in order to extract $n_{\pm}$. In fact, $n_{0}$ is the average refractive index or the geometric mean of $n_{x}=\sqrt{\epsilon_{xx}}$ and $n_{y}=\sqrt{\epsilon_{yy}}$, and $\theta$ represents the in-plane anisotropy, with $\theta=0$ being the isotropic case. We can then express in a different way the average of the in-plane dielectric functions as:
$\bar{\epsilon}=n_{0}^{2}\cosh{\theta}$ (25)
and the linear birefringence term as:
$\delta=n_{0}^{2}\sinh{\theta}$ (26)
and write Eq. 24 as:
$n^{2}_{\pm}=n_{0}^{2}\left(\cosh{\theta}\pm\sqrt{\sinh^{2}{\theta}-\alpha^{2}}\right)$ (27)
with $\alpha=\epsilon_{xy}/n_{0}^{2}$ being a dimensionless antisymmetric parameter and much smaller than 1 for typical magneto-optical experiments.
For the isotropic case (as in Ref. [22]) where $\epsilon_{xx}=\epsilon_{yy}=\epsilon$ so that $\delta=0$ Eq. 24 can be simplified to $n^{2}_{\pm}=\epsilon\pm\epsilon_{xy}$ and expanded to obtain:
$n_{\pm}=n_{0}\pm\frac{\epsilon_{xy}}{2n_{0}}+\mathcal{O}(\epsilon_{xy}^{2})$ (28)
with a linear in $\epsilon_{xy}$ term. The eigenmodes acquire a phase:
$\phi_{\pm}=k_{\pm}d=\frac{\omega n_{\pm}d}{c}$ (29)
through a slab of thickness $d$ where $c$ is the speed of light in vacuum and $k_{\pm}$ is the wavenumber for RH and LH circular polarization. Plugging in Eq. 28 and $\phi_{0}=\omega n_{0}d/c$ gives:
$\phi_{\pm}=\phi_{0}\pm\frac{\omega d}{2c}\frac{\epsilon_{xy}}{n_{0}}$ (30)
The complex Faraday angle is defined as:
$\Theta=\theta_{F}+i\eta=\frac{1}{2}\left(\phi_{+}-\phi_{-}\right)$ (31)
and therefore allow us to write the relation between the off-diagonal terms of the dielectric tensor to the Faraday angle:
$\epsilon_{xy}(\omega)=\frac{2c}{\omega d}n_{0}(\omega)\Theta(\omega)$ (32)
or the same for the conductivity tensor via:
$\sigma_{xy}(\omega)=-i\omega\varepsilon_{0}\epsilon_{xy}(\omega)$ (33)
This derivation highlights the direct relation between the linear in $\epsilon_{xy}$ term to the Faraday angle rotation and the fact that the off-diagonal terms are those responsible for it.
If the crystal is slightly birefringent so that $\delta\ll\bar{\epsilon}$ the square root in Eq. 24 can be expanded as:
$\sqrt{\delta^{2}-\epsilon_{xy}^{2}}=\delta-\frac{\epsilon_{xy}^{2}}{2\delta}+\mathcal{O}(\epsilon_{xy}^{4})$ (34)
which does not have any linear in $\epsilon_{xy}$ term. The Faraday angle rotation is obtained from measurements in positive and negative magnetic fields. Therefore, the contribution from the birefringent term will be canceled out since (i) the magnetic field does not affect $\delta$ directly, (ii) $\epsilon_{xy}$ is squared and symmetric for negative and positive fields. As a result, the birefringent term will not contribute to the total Faraday angle rotation. In particular, if $\delta\ll\epsilon_{xy}$ then the rotation is still dominated by $\epsilon_{xy}$ and the circular basis is a good approximation. However, if $\delta\gg\epsilon_{xy}$ linear birefringence dominates and the Faraday effect is strongly suppressed or masked, which is not our case, as can be seen in the data.